Find the integral surface of the linear PDE
x(y²+z)p - y(x²+z)q = (x²-y²)z , which contains the straight line x+y =0 , z=1.
1
Expert's answer
2019-08-27T12:00:10-0400

Auxiliary equations are

"\frac{dx}{x(y^2+z)}=\frac{dy}{-y(x^2+z)}=\frac{dz}{z(x^2-y^2)}"

"\frac{xdx+ydy-dz}{x^2(y^2+z)-y^2(x^2+z)-z(x^2-y^2)}=\frac{xdx+ydy-dz}{0}" (1)

"\frac{\frac{dx}{x}+\frac{dy}{y}+\frac{dz}{z}}{y^2+z-(x^2+z)+(x^2-y^2)}= 
\frac{\frac{dx}{x}+\frac{dy}{y}+\frac{dz}{z}}{0}" (2)

from (1)

"x^2+y^2-2z=C_1" (3)

from (2)

"log(xyz)=C_2,\ or \ xyz=C_3" (4)

parametric equation of straight line is

"(x=t,y=-t,z=1)" (5)

On substituting (5) in (3) and (4) we get

"2t^2-2=C_1,\ -t^2=C_3"

Eliminating t

"-2C_3-2=C_1,\ or F(C_1,C_3)=C_1+2C_3+2=0"

Therefor

"x^2+y^2-2z+2xyz+2=0" is integral surface, which contains the straight line.

Answer: x^2+y^2-2z+2xyz+2=0